OpenStudy (anonymous):

Solve the equation: 3 cot^2(x) - 1= 0

OpenStudy (amorfide):

you know that \[cotx=\frac{ 1 }{tanx }\] or you know that \[\cot^{2}x= \frac{ \cos^{2}x}{ \sin^{2}x }=\frac{ 1 }{ \tan^{2}x }\]

OpenStudy (amorfide):

you can rearrange your expression to get \[3\cot^{2}x=1\] \[\cot^{2}x=\frac{ 1 }{ 3 }\]

OpenStudy (amorfide):

this should help you enough

OpenStudy (anonymous):

yes i see how this works. just testing this out

OpenStudy (anonymous):

answer should be \[\sqrt{3}/3\]

OpenStudy (amorfide):

\[\frac{ 1 }{ \tan^{2}x }=\frac{ 1 }{ 3 }\] \[3=\tan^{2}x\] \[\pm \sqrt{3}=tanx\] \[x=60^{o},-60^{o}\]